random sample of 200 measurements give a mean of 50 and standard deviation of 9. at 5% level of significance, test the hypothesis that the population mean is 52.
Answer by stanbon(75887)

random sample of 200 measurements give a mean of 50 and standard deviation of 9. at 5% level of significance, test the hypothesis that the population mean is 52.
--------
Ho: u = 52 (claim)
Ha: u is not equal to 52
----
sample mean = 50
z(50) = (50-52)/[9/sqrt(200)] = -3.1427
----
p-value = P(z<-3.1427) = 0.00083705
-------
Since the p-value is less than 5%, reject Ho.
The test results do not support the claim.
================
Cheers,
Stan H.
================